Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2024 logbook, Contact: F.Wolff-Fabris, A.Galler, H.Sinn  Not logged in ELOG logo
Message ID: 178     Entry time: 26 Sep 2024, 23:12
Author: Chan Kim 
Group: SPB 
Subject: Shift summary 

X-ray delivery

  • 6 keV, ~2.7 mJ, stable delivery

Optical laser delivery

  • ns laser for particle visualization

Achievements/Observations

  • NKB focus optimized
  • ns laser aligned with particle beam
  • particle injector position optimized
  • measured 3 samples: roughly 2 hours per each sample
  • one sample was measured with and without He gas
  • Julabo chiller (AGIPD) is still on target (-32 degree)

Issues

  • Beam drift in vertical direction: Rz has to be corrected during the measurement and moved by 0.2 mm in total in whole day (~18 hours)
ELOG V3.1.4-7c3fd00